Suppose you are going to receive $11,000 per year for 6 years. The appropriate interest rate is 6 percent.

a. What is the present value of the payments if they are in the form of an ordinary annuity?

b. What is the present value if the payments are an annuity due
c. Suppose you plan to invest the payments for 6 years, what is the future value if the payments are an ordinary annuity?
d. Suppose you plan to invest the payments for 6 years, what is the future value if the payments are an annuity due?
Answer #1

a.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=11000[1-(1.06)^-6]/0.06

=11000*4.91732433

=$54090.57(Approx).

b.Present value of annuity due=Present value of annuity*(1+Rate)

=54090.57*1.06

=$57336.00(Approx).

c.Future value of annuity=Annuity[(1+rate)^time period-1]/rate

=11000[(1.06)^6-1]/0.06

=11000*6.97531854

=$76728.50(Approx).

d.Future value of annuity due=Future value of annuity*(1+rate)

=76728.50*1.06

=$81332.21(Approx).